[INSTRM-2065] Modify the port assignment for power control of the old E-box Created: 07/Sep/23 Updated: 24/Nov/23 Resolved: 24/Nov/23 |
|
| Status: | Done |
| Project: | Instrument control development |
| Component/s: | ics_pebActor |
| Affects Version/s: | None |
| Fix Version/s: | None |
| Type: | Task | Priority: | Normal |
| Reporter: | yuki.moritani | Assignee: | chyan |
| Resolution: | Done | Votes: | 0 |
| Labels: | near-term | ||
| Remaining Estimate: | Not Specified | ||
| Time Spent: | Not Specified | ||
| Original Estimate: | Not Specified | ||
| Description |
|
Following discussions at PFI telecon on 6-7 Sep 2023, the wiring of AG power supply will be modified so that one power supply covers odd/even AGs. The PDU post assignment will be changed accordingly [*], and modification will be done for only the old E-box so fa. ics_pebActor needs to be updated to handle power control for both the new and old E-box.
[*] AG2 and AG5 will be swapped. |
| Comments |
| Comment by yuki.moritani [ 17/Oct/23 ] |
|
Note: because the wiring of the AG3 and AG4 temperature sensor in the old E-box was modified again, software modified accordingly. |
| Comment by yuki.moritani [ 19/Oct/23 ] |
|
The modification for AG3 and AG4 will be revered as it turns out the wring is the same between the tow E-boxes. |
| Comment by yuki.moritani [ 24/Nov/23 ] |
|
The modification has been merged to master (and tagged as 1.1.5), so I close this ticket. If the problem happens, it will be addressed under a new ticket. |